Lindsay is building a mobile to hang over her baby's crib. She hangs a 0.20kg toy sailboat 0.01m from the left end and a 0.015kg toy truck 0.20m from the right end of a bar 0.50m long. If the lever arm itself has negligible mass, where must the support string be placed so that the arm balances?


1
Expert's answer
2020-12-10T11:03:33-0500
"m_1gr_1=m_2gr_2\\(0.2)(x-0.01)=(0.5-x-0.2)(0.015)\\x=0.03\ m\text{ from the left end}"